OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. HELPPPP ASAP

OpenStudy (anonymous):

what are the answer choices. y = 3x -1 should be the answer.

OpenStudy (anonymous):

y = m x + b 5 = 3*2 + b Solve for b b = -1 Put the parts together. y = 3 x + (-1) y = 3 x - 1 GT's answer is correct.

OpenStudy (anonymous):

His answer choices were given in a different form. Basically, one of the choices was: y - 5 = 3(x-2), which is the same thing as y = 3x-1.

OpenStudy (anonymous):

Sorry. I didn't have prior knowledge of external solution offerings.

OpenStudy (anonymous):

He posted the same question down below and I was working on it with him.